NaH2PO4 (1M) and Na2HPO4 (1M)

a. Calculate the pH. (show all the work)

b. Write the balance equilibrium.
Answer #1

a)

Henderson - Hasselbalch equation which give pH of a buffer solution is

pH = pKa + log([A-]/[HA])

where,

A- = conjucate base, HPO42-

HA = weak acid , H2PO4-

pKa of H2PO4- = 7.21

pH = 7.21 + log( 1M/1M)

pH = 7.21 + 0

pH = 7.21

b)

H2PO4- (aq) + H2O(l) <--------> HPO42-(aq) + H3O+(aq)